OpenStudy (anonymous):

Sailors need to consider the speed of the wind when adjusting the sails on their boat. The force F (in pounds per square foot) on a sail when the wind is blowing perpendicular to the sail can be modeled by the function F = 0.004v2 where v is the wind speed (in knots). Find the wind speed that will produce a force of 2.5 pounds per square foot on a sail.

OpenStudy (anonymous):

put F=2.5 2.5=0.004 v^2 2.5/0.004 v^2 625=v^2 v=25 or -25(rej.)

OpenStudy (anonymous):

thank you so much!!!!!!

OpenStudy (anonymous):

No problem.